OpenStudy (anonymous):

solve the equation then check your solution 7x+10=-18

OpenStudy (anonymous):

7x=-18-10 7x=-28 divide both sides by 7. x=-4 :)

OpenStudy (shayaan_mustafa):

Hi earsopen :) How are you dear? Welocme to Openstudy :D Now further, as you said to check the solution, Substitute x=-4 in the given question. 7(-4)+10=-18 -28+10=-18 -18=-18 You see both sides are equal. So its mean answer is correct.